# Reinforcement Learning for Control of Evolutionary and Ecological Processes

### v2

@@
Bryce Allen Bagley Mathematical Medicine Group, Stanford University Petritsch Laboratory, Department of Neurosurgery, Stanford Medical School Physician-Scientist Training Program, Stanford Medical School Correspondence: bbagley@stanford.edu Navin Khoshnan Mathematical Medicine Group, Stanford University Petritsch Laboratory, Department of Neurosurgery, Stanford Medical School Claudia K Petritsch Mathematical Medicine Group, Stanford University Petritsch Laboratory, Department of Neurosurgery, Stanford Medical School
@@
As Evolutionary Dynamics moves from the realm of theory into application, algorithms are needed to move beyond simple models. Yet few such methods exist in the literature. Ecological and physiological factors are known to be central to evolution in realistic contexts, but accounting for them generally renders problems intractable to existing methods. We introduce a formulation of evolutionary games which accounts for ecology and physiology by modeling both as computations and use this to analyze the problem of directed evolution via methods from Reinforcement Learning. This combination enables us to develop first-of-their-kind results on the algorithmic problem of learning to control an evolving population of cells. We prove a complexity bound on eco-evolutionary control in situations with limited prior knowledge of cellular physiology or ecology, give the first results on the most general version of the mathematical problem of directed evolution, and establish a new link between AI and biology.
@@ INTRODUCTION
One of the most productive areas in the study of evolutionary dynamics has been Evolutionary Game Theory (EGT).1, 2, 3, 4, 5, 6, 7 EGT combines elements of the discrete games from economics, the differential games primarily from engineering, and the mathematical and biological literature on evolution and ecology, and uses them to predict how evolutionary processes determine the frequencies of different types in a population. [^1] The term ”types” in EGT is a deliberately general one, as the mathematical tools can be used for the study of genotype frequencies in a population, foraging strategies in macro-organisms, cultural changes, and even collective behavior in response to economic policies.
A key goal of predicting evolutionary dynamics is to be able to intervene and steer evolution towards a desired state. This would represent a quantitative form of directed evolution, an approach of artificial selection already used on bacteria to optimize biomolecular catalysts for industrial manufacturing of pharmacologics, agricultural chemicals, and more.8, 9, 10, 11 This is essentially an approach of applying artificial selection in much the same fashion that natural selection optimizes organisms in nature. While the evolutionary procedures applied in these projects have been relatively simple in comparison with the complexity of evolution in nature, their existing success is indicative of the possibilities which lie in exploiting more of evolutionary – and especially eco-evolutionary – dynamics. Accomplishing this will require mathematical tools for guiding dynamics in a fine-grained manner, and such tools will themselves require the development of new theoretical and computational results. In this paper we expand the theoretical methods available for evolutionary dynamics via integrating contemporary research in control theory and the theory of artificial intelligence.
@@
In medical contexts one finds problems of directed evolution in the choice of therapies for tumors and other pathologies whose ability to evolve drug resistance is a key obstacle in curing many patients. There is been growing interest within clinical medicine regarding the possibility of applying EGT to the development and optimization of cancer therapies, immunology, and infectious disease treatment, but thus far there has been limited ability to integrate the enormous quantities of available biomedical data. Additionally, there are very few studies of treatment courses with meaningful dynamics over time, particularly because of the difficulty of experimenting with combination and time-course therapies. Instead, the current standard of care in cancer is generally to use one or two chemotherapeutics along with surgery and/or radiation. If the patient is not cured, the tumor develops resistance to those drugs, and a new drug or drug combination must be used. However, as discussed in 7, 3, 6, 5 and many others, there is growing evidence that therapies which vary over time in more strategic and complex ways have great potential to improve cancer outcomes. Similarly, bacterial and viral infections develop resistance to antibiotics and antivirals via evolutionary processes very much analogous to those in tumors.
12, 13 A recent resurgence of EGT in the social sciences14, 4 has seen analogous ideas return to substantial use in their intersections with engineering. This is most commonly seen in areas such as transit, energy distribution, computer usage, and other resource problems determined by the behaviors of large numbers of individuals belonging to a smaller number of types which each correspond to some form of pattern in a user’s problem-relevant behavior.15, 16, 17 Our work is distinct from this approach in that we are not using heuristic genetic algorithms for control, but rather are using rigorous methods for the control of evolution.
In economics, engineering, and biomedical research alike, here has been relatively little work applying contemporary advancements in control theory and related disciplines to questions of evolutionary systems. There is a recent discussion of the use of dynamic programming for the optimization of cancer therapy,18 however, we have been unable to find any such work beyond that paper. The massive advancements in the mathematical study of control systems and artificial intelligence from recent decades have not yet crossed over into the study of evolutionary dynamics. This represents a fertile ground for the use of new tools and the answering of previously unstudied – and in some cases perhaps even previously unposable – questions.
@@
Here we explore two key ideas. First, questions of how to prove a rigorous algorithmic bound on the computational problem of optimal control of evolution. Second, the duality in the nature of the underlying process as both a competition and a computation. The former has clear connections with what we discuss above about the history and present state of the cross-talk between game theory and evolutionary biology. [^2] This conceptual view is ubiquitous in theoretical biophysics, systems biology, and adjacent fields, and a recent review article offers a helpful explanation of them in terms of potential therapeutic applications.19 The latter, however, is perhaps a subtler one, and is related to a radical yet compelling proposal by Krakauer, Bertschinger, Olbrich, Flack, and Ay – theoretical biophysicists and theoretical biologists working in the space of complex systems science.20 They propose that individuality exists not as a binary property, but a spectral one with its origins in information theory. Instead of arbitrary human-imposed boundaries, one instead looks at how the mutual information of different partitionings varies and thus defines degrees of individuality.
We do not make use of the explicit mathematical description they offer, but take inspiration from its conceptual implications. Instead, we note that from a pure information perspective a change in a molecule’s structure is fundamentally no more or less an information transform than a molecule being moved from one partition to another. We can imagine attaching to any molecule a hypothetical string of bits to serve as an address of sorts. One value for the bit string could say that it is within a biological cell, another that it is outside the cell. This equivalence will be discussed further in Computational Duality and Multicules of the .
@@
Further, we make use of Osband and Van Roy’s 21, 22 work on model-based reinforcement learning (RL) of Markov Decision Processes, in which they proved the first unified regret bounds for model-based RL of quadratic systems – like EGT – which to our knowledge remain state of the art to this day. Linking these results from the theory of artificial intelligence with the specific function classes [^3] A term used in machine learning and artificial intelligence to denote a collection of functions with one or more relevant properties in common. of interest in EGT, we introduce and develop a mathematical formalism for an evolutionary control theory via a combination of EGT, reinforcement learning, and control systems. In addition to defining and exploring this formalism, we derive bounds on learning to control directed evolution which match the state-of-the-art bounds in the theory of artificial intelligence more broadly. However, much of the value of theory comes from eventual application to practice, and as discussed by Traulsen and Glynatsi there is a growing divide between theory and practice in EGT.23 Our goal in this paper is to develop a novel approach which introduces new techniques into the theoretical literature of EGT while focusing on its ultimate use in experiment and application.
@@
This general form of nonlinear dynamical system can thus be used to predict how the composition of the overall population will change over time. Extensions often account for mutations, which in the continuous case essentially transfer infinitesimal fractions of the population from one type to another, but the general principle holds.1, 2, 4, 24
@@
In the late 20th century (and especially the 21st thus far) quantitative biology largely split into theoretical biology and bioinformatics. Compared with the deeply interlinked nature of theory and data analysis in the first half of the 20th century, comparatively little interaction occurred between them in the latter half. Despite being theoretical, EGT has still provided useful insights even in abstract yet practical contexts. It offers explanations of why intuitively strange situations in ecology can remain stable, and in recent years has seen early application to analyzing tumor progression and microbial evolution of antibiotic resistance. However, these efforts have in some ways been constrained by the methods used. In most applications the number of types $q$ might be three or four, and the entries of $\mathbf{F}$ will be bird’s-eye-view estimates from large-scale observations of bacterial and cancer cell populations.
Bioinformatics tools which have emerged in recent decades may allow quantitative researchers to generate higher-fidelity approximations of $\mathbf{F}$ for $q$ large enough to make clinically relevant predictions beyond the simple test cases being assessed at present. Such work is critical, and a goal of this study is to begin the creation of a new class of tools for such researchers to use. In particular, work in the area of network inference algorithms has made it increasingly possible to estimate interaction effects within cells and between them. This represents an opportunity for a new approach to eco-evolutionary dynamics which can make use both of contemporary methods in biomedical data sciences and of rich theory.
@@ RESULTS
As discussed in the Methods subsection Computational Duality and Multicules, one of the new methods we introduce is to model multiple types of biophysical processes as all representing a single unifying sort of computation. For example, the conversion of one molecule into another by an enzymatic process is a computation as it changes the set of interactions the matter from the original molecule can engage in. Likewise, the movement of a molecule from a cell into the intercellular space – or vice versa – also changes the set of interactions that molecule can engage in. We introduce the term multicules to denote the units of this unifying computational representation.
@@
There are a few options in terms of how to model a system where data is available in the form of values for $c$ and limited data on $x$, while the dynamics themselves are in truth defined in terms of $x$. One option is to assume a given estimate of $c$ maps onto an estimated $x$ (and vice versa), and have the model be a hybrid of dynamics over $c$ and $x$. For the purposes of this study, the approach will be to assume a pseudo-invertable [^4] See Section 1 for a discussion of the validity of the assumption of pseudo-invertability. linear map $\mathbf{B}x=c$ which is time-invariant. Incorporating noise will be relevant later, so we will use the notation $\mathbf{\tilde{B}}=\mathbf{B}x+\xi$ with $\xi$ being a vector of zero-mean noise. For the time being, however, we will assume perfect ”estimates” and ignore the $\xi$ component of $\mathbf{\tilde{B}}$. Further, we define a vector of inputs to the ecology $u$, and a matrix $\mathbf{Z}$ mapping inputs to multicules.
However, it is not in fact required that we have any knowledge of $\mathbf{B}$, $\mathbf{G}$, or any other parameter matrices a priori. This is where the application of reinforcement learning will become central to later sections of the study. For now, we must first define a model of the eco-evolutionary process. The rest of this section will be a discussion of the model and how we can convert it into a Partially Observable Markov Decision Process (POMDP) so as to enable the use of the extensive research in control systems engineering and the theory of artificial intelligence.
0.1 Dynamical System Model for Multicule Eco-Evolutionary Dynamics
@@
The underlying mathematical model is described entirely with respect to multicule ”computations” on $x$ as
@@
With $\mathbf{C}_{B}^{T}$ being a matrix whose diagonals are the entries of $\mathbf{B}x$. In other words, $\mathbf{C}_{B}$ is defined as
@@ 0.2 Non-static Ecology
It is important to note that the form of $\mathbf{\Delta}\mathbf{U}\mathbf{V}$ captures modifications to the ecology of an eco-evolutionary system which are separate from the direct multicules introduced via an external input into the system. The matrices $\mathbf{\Delta}$ and $\mathbf{V}$ are needed to reshape the dynamics of $\mathbf{U}$, and the three matrices in concert describe the effects of inputs on the ecological state of the model just as $\mathbf{Z}$ describes the inputs’ direct mapping to $x$ where relevant.
@@
If we proceed through the matrix multiplications as discussed in Section 1, we derive the completed model of the evolutionary dynamics with the estimation procedures taken into account.
@@
Note that $\mathbf{Z}$ is equivalent to $\mathbf{B}^{-1}\tilde{\mathbf{Z}}$, to streamline notation. Entries of $\mathbf{Z}$ serve to map external inputs into the multicules they become within the system, in this study according to the functional form of $u$ as decays from a given input point. In terms of pharmaceuticals, the term including $\mathbf{U}$ can be thought of as the effects of active drug molecules, and $\mathbf{Z}u$ as combining the rates of excretion and the rates of metabolism into pharmacologically inactive compounds.
0.3 POMDP for Multicule Eco-evolutionary Models
This is a nonlinear dynamical system just as with Equation 1, and in the next section we proceed to a discussion of how model predictive control can be used to influence it. Finally, for the sake of our later analysis, it can be useful to convert the system in Equation 8 into a single, equivalent higher-order equation. The derivation is somewhat messy, so they can be found in Section 1 as well. Recalling our earlier mention of a ”bookkeeping” variable, we note that it is not illustrative in the context of the model’s dynamics, but does verify the model’s property of being a quadratic dynamical system. That property will be important later, as regret bounds for model-predictive reinforcement learning are not known for all but a small number of very specific equations and a much smaller number of model classes – linear, quadratic, and generalized linear models being the only such general classes for which there are to our knowledge adequate bounds.
@@ 0.3 POMDP for Multicule Eco-evolutionary Models
As we will soon show, the use of reinforcement learning means that there are minimal requirements of a-priori knowledge of a system. Beyond specifying definitions of multicules and input compounds, we can leave all parameter matrices unknown at the start of the problem. this will of course increase the number of trials needed to learn a model when compared with learning given partial prior knowledge of parameters, but the worst-case bounds described in LABEL:accuracy_section are valid for the cases of any amount of prior knowledge, including the case of no such knowledge. Armed with this single-equation POMDP, we proceed to develop the optimal control results.
0.4 Parameter Learning
@@
One will inevitably need to select a subset of variables to create a Reduced-Order Model (ROM) for the POMDP $\Psi$, given the enormous dimensionality of the ideal approximation: a Full-Order Model incorporating all multicules involved in the system. There will be a lower bound on the error resulting from the difference between the ROM and the FOM, expressed as the residual, which is the difference between the state resulting from a ROM and that which would be seen for the same variable subset in an FOM. We denote the magnitude of this error as $\varepsilon_{O}$. A number of authors have described error bounds for various types of dynamical systems, and this continues to be an area of active research.25, 26, 27, 28, 29, 30 To our knowledge the best bound on this error for nonlinear dynamical systems is given by Chellappa et al in 31. Their method is rather technical, and translating their error bounds to the context of our problem in detail is beyond the scope of this paper. We note the existence of such methods both to acknowledge an interesting and highly relevant body of literature, and to indicate the direction one could pursue in proving detailed bounds on $\varepsilon_{O}$. In addition, there is some fundamental error limit $\varepsilon_{\text{min}}$ below which a quadratic model for a given biological system could not be further optimized, regardless of whether or not it were reduced-order. Thus the problem we analyze here is more precisely one of reinforcement learning to minimize any error beyond the minimum set by those two limits, and optimal treatment selection given such an optimized model.
The error bounds for reinforcement learning of POMDPs proven across 32, 22, 21 are both state of the art [^5] Per our literature review and personal communication with B. Van Roy and M. Kochenderfer. and highly general compared with most results in the field, which makes them very suitable for this problem. A lengthier discussion of these bounds in our case can be found in Supplemental Information Section 4, and relate specifically to the algorithm Posterior Sampling Reinforcement Learning (PSRL).22 In summary, we use their results to derive regret for using the PSRL algorithm to both learn the parameter matrices and select inputs to optimize trajectories through evolutionary space.

### v3

@@
Bryce Allen Bagley Mathematical Medicine Group, Stanford University, Stanford, CA 94305, USA Department of Neurosurgery, Stanford Medical School, Stanford, CA 94305, USA Physician-Scientist Training Program, Stanford Medical School, Stanford, CA 94305, USA Correspondence: bbagley@stanford.edu Navin Khoshnan Mathematical Medicine Group, Stanford University, Stanford, CA 94305, USA Department of Neurosurgery, Stanford Medical School, Stanford, CA 94305, USA Claudia K Petritsch Mathematical Medicine Group, Stanford University, Stanford, CA 94305, USA Department of Neurosurgery, Stanford Medical School, Stanford, CA 94305, USA Maternal & Child Health Research Institute, Stanford University School of Medicine, Stanford, CA 94305, USA Stanford Cancer Institute, Stanford University School of Medicine, Stanford, CA 94305, USA
@@
As research in evolutionary dynamics and Evolutionary Game Theory moves from the realm of theory into application in oncology, algorithms that account for physiologic variation are needed to move beyond simple models. Yet few algorithms exist that integrate these ecological and physiological factors which are central to evolution in biological systems. We introduce a new model of evolutionary games which accounts for ecology and physiology by framing both as computations. Leveraging this new model, we develop an algorithmic approach for controlling evolution via methods from Reinforcement Learning, a branch of Machine Learning, and prove rigorous algorithmic bounds on the complexity of solving this problem. This approach enables us to obtain first-of-their-kind results on the algorithmic problem of learning to control an evolving population of cells. We prove a complexity bound on controlling evolution in situations with limited prior knowledge of cellular physiology or ecology, give the first results on the most general version of the mathematical problem of directed evolution, and establish a new link between AI and biology.
@@ INTRODUCTION
In medical contexts, the clearest application of Evolutionary Game Theory (EGT).1, 2, 3, 4, 5, 6, 7 is to modeling the physiology of cancers and their responses to therapies with the goal to optimize cancer treatment.8, 9, 7, 10, 11 There is growing interest within clinical medicine regarding the possibility of applying EGT to the development and optimization of cancer therapies, immunology, and infectious disease treatment, but thus far there has been limited ability to integrate the enormous quantities of available biomedical data. In addition, there are very few studies of treatment courses with meaningful dynamics over time, particularly because of the difficulty of experimenting with combination and time-course therapies. Instead, the current standard of care in cancer is generally to use one or two chemotherapeutics along with surgery and/or radiation. If the patient is not cured, the cancer develops resistance to those drugs through multi-faceted mechanisms, and a new drug or drug combination must be used. Mechanisms of resistance include, among others, modulation of the immune system, changes to non-coding regions which interfere with therapy efficacy, and mutations and epigenetic changes which alter features of cancer cells such as signalling pathways, drug transport into and out of cells, target site structure, and drug metabolism.12 However, as discussed in 7, 3, 6, 5 and many others, there is growing evidence that therapies which vary over time in more strategic and complex ways have great potential to improve cancer outcomes. Similarly, bacterial and viral infections develop resistance to antibiotics and antivirals via evolutionary processes very much analogous to those in cancers.
EGT uses combined elements of the discrete games from economics, the differential games primarily from engineering, and the mathematical and biological literature on evolution and ecology to predict how evolutionary processes determine the frequencies of different types in a population. [^1] The term ”types” in EGT is a deliberately general one, as the mathematical tools can be used for the study of genotype frequencies in a population, foraging strategies in macro-organisms, cultural changes, and even collective behavior in response to economic policies. In biomedical contexts, there has been particular interest in the application of EGT to the study of cancer and therapy resistance,3, 5, 6, 7, 13, 14, 15 and this is likewise our focus in this paper.
A key goal of predicting evolutionary dynamics is to be able to intervene and steer evolution towards a desired state. This would represent a quantitative form of directed evolution, an approach of artificial selection already used on bacteria to optimize biomolecular catalysts for industrial manufacturing of pharmacologics, agricultural chemicals and more.16, 17, 18, 19 This is essentially an approach of applying artificial selection in much the same fashion that natural selection optimizes organisms in nature, and the existing experimental success is indicative of the possibilities which may lie in more quantitatively-driven use. Accomplishing this will require mathematical tools for guiding dynamics in a fine-grained manner, and such tools will themselves require the development of new theoretical and computational results.
@@
Here we explore two key ideas. First, questions of how to prove a rigorous algorithmic bound on the computational problem of optimal control of evolution. Second, the duality in the nature of the underlying process as both a competition and a computation. The former has clear connections with what we discuss above about the history and present state of the cross-talk between game theory and evolutionary biology. [^2] This conceptual view is ubiquitous in theoretical biophysics, systems biology, and adjacent fields, and a recent review article offers a helpful explanation of them in terms of potential therapeutic applications.20 The latter, however, is perhaps a subtler one, and is related to a radical yet compelling proposal by Krakauer, Bertschinger, Olbrich, Flack, and Ay.21 They propose that individuality exists not as a binary property, but a spectral one with its origins in information theory. Instead of arbitrary human-imposed boundaries, one instead looks at how the mutual information of different partitionings varies and thus defines degrees of individuality. This will be discussed further in the Methods section on Computational Duality.
@@
Our goal in this paper is to develop a novel approach which introduces new techniques into the theoretical literature of EGT while focusing on its ultimate use in experiment and application. To accomplish this, we make use of Osband and Van Roy’s work on model-based reinforcement learning (RL) of Markov Decision Processes, in which they proved the first unified regret bounds for model-based RL of quadratic systems – like EGT – which to our knowledge remain state of the art to this day.22, 23 Linking these results from the theory of artificial intelligence with the specific function classes [^3] A term used in machine learning and artificial intelligence to denote a collection of functions with one or more relevant properties in common. of interest in EGT, we introduce and develop a mathematical formalism for an evolutionary control theory via a combination of EGT, reinforcement learning, and control systems. In addition to defining and exploring this formalism, we derive bounds on learning to control directed evolution which match the state-of-the-art bounds in the theory of artificial intelligence more broadly. However, much of the value of theory comes from eventual application to practice, and as discussed by Traulsen and Glynatsi there is a growing divide between theory and practice in EGT.24
@@
This general form of nonlinear dynamical system can thus be used to predict how the composition of the overall population will change over time. Extensions often account for mutations, which in the continuous case essentially transfer infinitesimal fractions of the population from one type to another, but the general principle holds.1, 2, 4, 25
@@
Compared with the deeply interlinked nature of theory and data analysis within bioscience in the first half of the 20th century, comparatively little interaction occurred between them in the latter half. Despite being theoretical, EGT has still provided useful insights even in abstract yet practical contexts. It offers explanations of why intuitively strange situations in ecology can remain stable, and in recent years has seen early application to analyzing cancer progression and microbial evolution of antibiotic resistance. However, these efforts have in some ways been constrained by the methods used. In most applications the number of types $q$ might be three or four, and the entries of $\mathbf{F}$ will be bird’s-eye-view estimates from large-scale observations of bacterial and cancer cell populations.
Bioinformatics tools which have emerged in recent decades may allow quantitative researchers to generate higher-fidelity approximations of $\mathbf{F}$ for $q$ large enough to make clinically relevant predictions beyond the simple test cases being assessed at present. Such work is critical, and a goal of this study is to begin the creation of a new class of tools for such researchers to use. In particular, work in the area of network inference algorithms has made it increasingly possible to estimate interaction effects within cells and between them. This represents an opportunity for a new approach to evolutionary dynamics which can make use both of contemporary methods in biomedical data sciences and of rich theory.
@@ RESULTS
Modeling the biophysics of disease physiology across multiple scales and diverse types of transport and biochemical processes is a significant mathematical challenge. As discussed in the Methods subsection on Computational Duality, one of the new methods we introduce is to model multiple types of biophysical processes as all representing a single unifying sort of computation. For example, the conversion of one molecule into another by an enzymatic process is a computation as it changes the set of interactions the matter from the original molecule can engage in. Likewise, the movement of a molecule from a cell into the intercellular space – or vice versa – also changes the set of interactions that molecule can engage in. We introduce the term multicules to denote the units of this unifying computational representation.
@@
Additionally, for interfacing with experimental research the model must be applicable to systems where data for values of $c$ is readily obtained while data for values of $x$ is more limited, even though the dynamics themselves are in truth defined in terms of $x$. One option is to assume a given estimate of $c$ maps onto an estimated $x$ (and vice versa), and have the model be a hybrid of dynamics over $c$ and $x$. For the purposes of this study, the approach will be to assume a pseudo-invertable [^4] See Supplemental Information Section 1 for a discussion of the validity of the assumption of pseudo-invertability. linear map $\mathbf{B}x=c$ which is time-invariant. Incorporating noise will be relevant later, so we will use the notation $\mathbf{\tilde{B}}=\mathbf{B}x+\xi$ with $\xi$ being a vector of zero-mean noise. For the time being, however, we will assume perfect ”estimates” and ignore the $\xi$ component of $\mathbf{\tilde{B}}$. Further, we define a vector of inputs to the ecology $u$, and a matrix $\mathbf{Z}$ mapping inputs to multicules. However, it is not in fact required that we have any knowledge of $\mathbf{B}$, $\mathbf{G}$, or any other parameter matrices a priori. This is where the application of reinforcement learning will become central to later sections of the study. For now, we must first define a model of the evolutionary process. The rest of this section will be a discussion of the model and how we can convert it into a Partially Observable Markov Decision Process (POMDP) so as to enable the use of the extensive research in control systems engineering and the theory of artificial intelligence.
One would ideally like there to be a fairly straightforward means of using existing experimental data to determine an initial model which the RL algorithm will then optimize. In what follows, we derive a model whose complexity can appear unreasonable. Indeed, our learning algorithm would produce a significantly underspecified set of matrices if it alone were used for inferring the model. The added complexity, however, is intentional and strategic for two reasons. First, it allows more direct input of known experimental values, which serve as a Bayesian prior for the model to be learned. This includes data on enzymatic reactions, estimates of the relationship between concentrations of different multicules, population fractions of different cell types, and more. Second, one can use the known values as an approximate ground truth for specifying various parameters, reducing – potentially quite substantially – the number of parameters to be derived from the model inferred via Reinforcement Learning. While this second approach may not be considered important in some use cases, it could be useful for many applications. This combination of benefits from the more mechanistic model justifies the complexity, and if a fully mechanistic model is not desired then one can simply use the full mathematical formulation for creating a sort of experiment-driven Bayesian prior over the space of possible models.
Dynamical System Model for Multicule Evolutionary Dynamics
@@
To begin, we want to ensure the underlying mathematical model is described entirely with respect to multicule ”computations” on $x$. As derived in Section 1 of the Supplemental Information, converting the standard EGT model to one based on multicules yields the following model.
@@
$\mathbf{C}_{B}^{T}$ is a matrix whose diagonals are the entries of $\mathbf{B}x$. In other words, $\mathbf{C}_{B}$ is defined as
@@
It is important to note that the form of $\mathbf{\Delta}\mathbf{U}\mathbf{V}$ captures modifications to the ecology of an evolutionary system which are separate from the direct multicules introduced via an external input into the system. The matrices $\mathbf{\Delta}$ and $\mathbf{V}$ are needed to reshape the dynamics of $\mathbf{U}$, and the three matrices in concert describe the effects of inputs on the ecological state of the model just as $\mathbf{Z}$ describes the inputs’ direct mapping to $x$ where relevant.
@@
If we proceed through the matrix multiplications as discussed in Supplemental Information Section 1, we derive the completed model of the evolutionary dynamics with the estimation procedures taken into account.
@@
Note that $\mathbf{Z}$ is equivalent to $\mathbf{B}^{-1}\tilde{\mathbf{Z}}$, to streamline notation. Entries of $\mathbf{Z}$ serve to map external inputs into the multicules they become within the system, in this study according to the functional form of $u$ as decays from a given input point. In terms of pharmaceuticals, the term including $\mathbf{U}$ can be thought of as the effects of active drug molecules, and $\mathbf{Z}u$ as combining the rates of excretion and the rates of metabolism into pharmacologically inactive compounds. As discussed at the beginning of the Results section, this complex system of equations enables a Bayesian prior hypothesis over the space of possible models and for the purposes of learning can readily reduce to a simpler model, which – given sufficient experimental data – can be re-expanded used as a mechanistic Bayesian posterior distribution over the space of possible models. This simpler model will be summarized using $\mathbf{A}$ to denote the largest matrix resulting from performing all the matrix multiplications in Equation 8.
POMDP for Multicule Evolutionary Models
Finally, for the sake of our later analysis, it can be useful to convert the system in Equation 8 into a single, equivalent higher-order equation. The derivation is somewhat lengthy, so it can be found in Section 1 of the Supplemental Information. The ”bookkeeping variable” mentioned earlier enables us to preserve the standard property of EGT models being quadratic dynamical systems. That property will be important later, as regret bounds for model-predictive reinforcement learning are not known for all but a small number of very specific equations and a much smaller number of model classes – linear, quadratic, and generalized linear models being the only such general classes for which there are to our knowledge adequate bounds.
@@
with $\vec{\xi}$ a zero-mean sub-Gaussian noise vector to account for model error (and absorbing the earlier $\xi$ term), multiplied by $(\delta t)$ so that it is proportional to the time-step length. There is thus a parameter $\sigma_{\psi}$ describing the maximum variance across entries of this sub-Gaussian noise vector. The noise is sub-Gaussian rather than Gaussian to prevent the mathematical difficulties connected with the MDP RL analysis on an unbounded range of outcomes, and this is also physically correct given conservation of mass. Because we can only observe a subset of the system’s components at any given time, this MDP is termed ”Partially Observable” and is thus a POMDP. Armed with this single-equation POMDP, we proceed to develop the optimal control results.
Parameter Learning
@@
One will inevitably need to select a subset of variables to create a Reduced-Order Model (ROM) for the POMDP $\Psi$, given the enormous dimensionality of the ideal approximation: a Full-Order Model incorporating all multicules involved in the system. There will be a lower bound on the error resulting from the difference between the ROM and the FOM, expressed as the residual, which is the difference between the state resulting from a ROM and that which would be seen for the same variable subset in an FOM. We denote the magnitude of this error as $\varepsilon_{O}$. A number of authors have described error bounds for various types of dynamical systems, and this continues to be an area of active research.26, 27, 28, 29, 30, 31 To our knowledge the best bound on this error for nonlinear dynamical systems is given by Chellappa et al in 32. Their method is rather technical, and translating their error bounds to the context of our problem in detail is beyond the scope of this paper. We note the existence of such methods both to acknowledge an interesting and highly relevant body of literature, and to indicate the direction one could pursue in proving detailed bounds on $\varepsilon_{O}$. In addition, there is some fundamental error limit $\varepsilon_{\text{min}}$ below which a quadratic model for a given biological system could not be further optimized, regardless of whether or not it were reduced-order. Thus the problem we analyze here is more precisely one of reinforcement learning to minimize any error beyond the minimum set by those two limits, and optimal treatment selection given such an optimized model.
The error bounds for reinforcement learning of POMDPs proven across 33, 23, 22 are both state of the art [^5] Per our literature review and personal communication with B. Van Roy and M. Kochenderfer. and highly general compared with most results in the field, which makes them very suitable for this problem. A lengthier discussion of these bounds in our case can be found in Supplemental Information Section 4, and relate specifically to the algorithm Posterior Sampling Reinforcement Learning (PSRL).23 In summary, we use their results to derive regret for using the PSRL algorithm to both learn the parameter matrices and select inputs to optimize trajectories through evolutionary space.